if x is inversely proportional to y, and x=60 when y=0.5, how do i find x when y=12?
Answer by Alan3354(69443)

if x is inversely proportional to y, and x=60 when y=0.5, how do i find x when y=12?
---------------
x inversely proportional --> x = k/y, some constant k over y.
x = k/y
xy = k
k = 60*0.5 = 30
--------------
@ y = 12: x = 30/12
x = 2.5